OpenStudy (anonymous):

Evaluate a-2 for a = 8.

OpenStudy (anonymous):

substitute it in. (8)-2=8

OpenStudy (anonymous):

order of operations next, try isolating the variable. What is the inverse operation of subtraction?

OpenStudy (anonymous):

addition

OpenStudy (anonymous):

Wait a min... a can not be 8...

OpenStudy (anonymous):

now im confused at what his saying XD

OpenStudy (anonymous):

yeah its a-8 a=8 so 8-2=?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

TADA!

OpenStudy (anonymous):

your welcome sincerely from the happy cow :D

OpenStudy (anonymous):

no your wrong

OpenStudy (anonymous):

o my D:

OpenStudy (anonymous):

could you explain why?

OpenStudy (anonymous):

1/-16 ,1/-64,1/64,-64

OpenStudy (anonymous):

these are the chocies

OpenStudy (anonymous):

i'm sorry idk the knowledge of a cow is limited :C

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@KissMyAxe can you finish helping me

OpenStudy (anonymous):

Still thinking @diannemarieee

OpenStudy (anonymous):

Any other information given??

OpenStudy (anonymous):

nope just that

OpenStudy (anonymous):

Of which topic is this question from

OpenStudy (anonymous):

its pre-algebra flvs

OpenStudy (anonymous):

is the question like this 'a^-2' because then u get 1/64

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

OOOO

OpenStudy (anonymous):

lol i thought it was subtraction!

OpenStudy (anonymous):

i was like "how is this possible!"

OpenStudy (anonymous):

U typed it wrong than @diannemarieee Nice of u to give the answers!

OpenStudy (anonymous):

@chucho78 Its AXE

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

@KissMyAxe hmm?

OpenStudy (anonymous):

Hey don't take it seriously but u your profile pic looks scary @diannemarieee